PA 20-7 You are given the following customer responses... You are given the following customer responses to a purchase intent survey for a new electric bicycle that has all its batteries built into the frame and hence almost looks exactly like a regular bike: Definitely would use: 12 Probably would use: 22 Might or might not use: 31 Probably would not use: 111 Definitely would not use: 201 Out of a target market of 1,500,000 potential customers, how many do you expect to adopt the product, assuming that you can reach a 5% awareness in the population? Assume Cprobably-0.2 and Cdefinite 0.4. Select the closest option. O 150 O 243 O 1830 O 24290 O 1500000PA 20-7 You are given the following customer responses to a purchase intent survey for a new electric bicycle...
Answer #1

As per results provided total people who took survey were 12+22+31+111+201 = 377

Potential customers are 1,500,000. Since awareness reached to 5% customer only, so potential customers who know about product will be 5% of 1,500,000 which is 75,000

Based on survey % wise response of the cutomers was

Definitely would use = 12/377 = 3.18%

Probably would use = 22/377 = 5.84%

Might or moght not use = 31/377 = 8.22%

Probably would not use = 111/377 = 29.44%

Definitely would not use = 201/377= 53.32%

So as per result expected people to use product are 3.18%(definite users) and 5.84% (Probable users)

Also as given probability of definite users using product is 0.4 and probability of probable users using product is 0.2

So total people who will adopt product will be 75,000* 3.18% *0.4 + 75,000*5.84%*0.2 = 954.90+ 875.33 =1830

c) option is correct

Thanks